Columbia Products produced and sold 1,300 units of the company’s only product in March. You have collected the following information from the accounting records:
  

Sales price (per unit) $ 125
Manufacturing costs:
Fixed overhead (for the month) 15,600
Direct labor (per unit) 7
Direct materials (per unit) 33
Variable overhead (per unit) 22
Marketing and administrative costs:
Fixed costs (for the month) 20,800
Variable costs (per unit) 3

  

Required:

a. Compute the following:

Variable Manufacturing cost per unit?

Full cost per unit?

Variable Cost per unit?

Full Absorption Cost per unit?

Prime Cost per unit?

Conversion cost per unit?

Profit margin per unit?

Contribution margin per unit?

Gross Margin per unit?
Answer #1

Variable manufacturing cost per unit = Direct material + direct labor + variable overhead

= 33+7+22

= $62 per unit

Full cost per unit = 7+33+22+3+(15600+20800)/1300

= $93 per unit

Variable cost per unit = 7+33+22+3 = $65 per unit

Full Absorption cost = Variable manufacturing cost + fixed overhead per unit

= 62+15600/1300

= $74 per unit

Prime cost = Material + labor

= 7+33

= $40 per unit

Conversion cost per unit = Labor cost + manufacturing overhead

= 7+22+15600/1300

= $41 per unit

Profit margin per unit = 125-(7+33+22+3)+(15600+20800)/1300

= $32 per unit

CM per unit = Sales price – variable costs

= 125 – 65

= $60 per unit

Gross margin = Sales – cost of goods sold

= 125-74

= $51 per unit